Following experimental values:

Room temp: 295K

Barometric pressure: 762.5mmHg

vapor pressure: 19.8mmHg

Volume O2 collected: 62mL

Density of H2O2: 1.01g/mL

% composition of H2O2: 3.02%

Volume H202: 5mL

Letter of unknown solution of H2O2: c

Volume of O2 collected: 56mL

1. Calculate the volume of O2 at STP.

2. Based on the reaction stoichiometry, calculate the number of moles of O2

3. Calculate the mass of O2:

4. Calculate molar volume at STP

5. Calculate the % error of the molar volume at STP

6.Calculate R in the ideal gas law.

7.Calculate the % error of the value of R.

8.Calculate the % composition of H2O2 in the unknown solution. (Round your answer to the nearest integer)
Answer #1

1+20, — 420 + 10, ( C, 62ure, r = 29ьк Su d=1.01g/me 20 = 14-8ш ми = 762.5 или 16 ham : 5 xol 3 у = 762-5 - 14 •g - 74 2-7 ши